You are given five values of mass and five values
Answer #1

Momentum = mass*velocity

So, Velocity = Momentum/mass

For this problem, write Momentums in decreasing order (descending)and write mass in incresing order(ascending)

like, Momentum (in kg m/s) = [2.2x104 , 1230 , 195 , 36.3 , 5.23 ]

And Mass (in kg) = [2.35 , 31.5 ,1.60x102 , 933 ,11800]

Largest possible speed = 2.2x104/2.35 = 9361.702128 m/s = 9361.7 m/s (approx)

Second largest speed = 1230/31.5 = 39.04761905 m/s = 39 m/s (approx)

Third largest speed = 195/160 = 1.21875 m/s = 1.2 m/s (approx)

Fourth largest speed = 36.3/933 = 0.038906752 m/s = 0.0389 m/s (approx)

smallest possible speed = 5.23/11800 = 4.43220339x10-4 m/s = 4.4x10-4 m/s (approx)